Based on the growth data of a bacterial culture presented below, determine μ (specific growth rate)....

  1. Based on the growth data of a bacterial culture presented below, determine μ (specific growth rate).

Time (Hour)

X (Cell Concentration, Cells/L)

1

1,100,000

2

1,600,000

3

2,620,000

4

3,650,000

5

6,080,000

6

9,900,000

7

18,000,000

8

28,100,000

9

39,100,000

10

70,100,000

11

102,000,000
Answer #1

Constant of proportionality = u (mu)

This is an index of growth rate.

u = [InNt - InNo] ÷ [t - to]

In is integral of.

Integral is calculated by finding log to the base 10 value multiplied by 2.303

[(log10 Nt - log10 No) X 2.303] ÷ ( t - to )

(t - to) = 11 - 1 = 10

Nt log to the base ten (logNt10) of 102000000 is = 8.01 ( from log table or web calculator)

logNo 10 of 1100000 is = 6.04 ( from log table or web calculator)

8.01- 6.04 = 1.97

1.97 X 2.303 = 4.54

So [(log10Nt - log10 No)2.303] ÷ (t-to)

Giving values = 4.54 ÷ 10

= 0.454/hr